OpenStudy (anonymous):

Prove that △ABC ∼ △DEC if AB is parallel to DE. file:///C:/Documents%20and%20Settings/name/My%20Documents/IMG_20140613_230938.jpg

OpenStudy (shamim):

figure plz

OpenStudy (neer2890):

upload your image..?

OpenStudy (anonymous):

Here's the image

OpenStudy (neer2890):



OpenStudy (neer2890):

HERE, ANGLE ABC= ANGLE DEC ANGLE ACB= ANGLE DCE Also, sum of all angles in a triangle is 180 so ANGLE BAC= ANGLE EDC

OpenStudy (neer2890):

ALSO, \[\frac{ AB }{ DE }=\frac{ BC }{ EC }=\frac{ AC }{ DC }\][ BY AAA PROPERTY] \[\Delta ABC \approx \Delta DEC\]